Service Observing Night UT20130517

Summary

  • Conditions never really met Sommariva constraints
  • 1x1200s Mauerhan SN2005bf
  • GRB 130427A 5x900s
  • Imaging on sn2010mc
  • Bias 3K, 8K done, other calibrations still needed

MODS

Files: mods1*.20130517.NNNN

Thin patchy clouds

03:06 Feige 66

acquire: R: 0001, confirm 0002, -0.37Xoff manually, 0003 re-confirm (still slightly offset)
dual grating R: 0004-0006 B: 0001-0003 blue grating B: 0004-0006

03:57 Sommariva N1625 (blue only) acquire: B: 0007-0008, confirm 0009 Seeing 1.4 on WFS. EL=39 (Airm=1.59) Stars 1 and 4 are a little off to the LL (+3 pix in X and Y) realign on frame 0009 using stars 1 and 4 only final position looks good. DIMM showing 0.9" at EL=61 but we are seeing ~1.3-1.4" on MODS at lower EL. Seeing exceeds Sommariva threshold.

04:18 Mauerhan SN2005bf (dual grating) acquire: R: 0007-0008 0008 stopped halfway through because 0007 was saturated on the ref star. 0008 still sat'd. confirm: 0009, stopped after 15 seconds.

Ref star spectrum: B: 0010 R: 0011 SN spectrum: B:0011 R:0012

05:00 back to Sommariva N1625 (blue only) acquire: 0013 0014, just apply sum of offsets from previous acq and confirm 0015 (small offset remaining) rerun modsAlign on 0013 0015 offsetpointing -0.0276 0.117 0.148 detxy rel re-confirm: 0016 alignment good, but tails visible on the stars(s)? remove slit and re-image: 0017 seeing still ~1.4, above limit for this project sources look a little elongated. Clear AO and allow to recollimate. 0018, still somewhat elongated but better than 0017. Gradient in FWHM across the image? Seeing is still bad. Discuss options 0019 Seeing still above Sommariva limit, moving on.

06:03 ND_grb acquire at PA=45 (grb_pa045.acq) R:0012 0013 confirm 0014 looks good grb.obs B: 0020-0025 R:0015-0020 last exposure interrupted. EL~27, seeing getting worse (1.5 on AGw)

08:11 back to Sommariva N1625 (blue only) SX M1 panic'ed, probably from low EL?

08:21 try again...collimation diverging. DGH to try a different field, clear active optics.

08:33 Collimate on nearby field...much better start, converging acquire 0028 0029 confirm 0030, added +0.32X +0.44Y manually 0031 confirm again, looks OK (stars 1 and 4) FWHM~1.1 arcsec.

08:59 n1625_1.obs started. EL~76 B: 0032- R: 0021- Had to abort just after the first exposure finished. Winds at limit, trending higher.

While closed: bias3K, bias8K

10:18 Re-opening

10:32 sn2010mc Dual imaging B: 0043- R: 0031- Significant DIQ changes across R image 0031, lesser on B0043 (but also worse seeing in the blue so perhaps it is being masked) Seeing not great, but twilight approaching and this was all I had that fit in the remaining time.

11:16 Closing dome, done for the night.

MODS is asleep.

-- DavidThompson - 17 May 2013
Topic revision: r3 - 17 May 2013, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback